A 22 g chunk of dry ice is placed in a empty 600
mL tightly closed vessel at 25 C . What would
be the final pressure inside the vessel if all CO2
gets evaporated?

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

w = 22g dry CO2 which will evaporate to develop P

m = 44

V = 0.6 litre

P = ?

T = 298 K

PV = wmRT

P×0.6=2044×0.0821×298


Pressure inside bottle = P + atm. pressure

Regards


Vikas TU
4 years ago
w = 22g dry CO2 which will evaporate to develop P
m = 44
V = 0.75 litre
P = ?
T = 298 K
PV = wmRT
P×0.6=2044×0.0821×298
⇒ P = 8.334 atm
Pressure inside bottle = P + atm. pressure
⇒ 8.334 + 1 = 9.334 atm